The cost in dollars of producing x headphones is C(x)= 1600+100x-0.1x2


a) calculate average cost of producing 200 headphones


b) calculate the marginal cost of producing the 200th headphone


1
Expert's answer
2021-03-31T10:12:29-0400

a). We allow for the number of headphones to take positive real values. The mean value (average cost) is: "\mu=\frac{C(x)}{x}=\frac{1600}{x}+1000-0.1x". We subsitute "x=200" and get: "\mu=8+1000-20=988"

b). The marginal cost is given by formula "C'(x)=100-0.2x". For "x=200" we receive: "C(200)=60"